# Entropic (Gromov) Wasserstein Flow Matching with GENOT

### v3

@@ 3.3 Combining Linear and Quadratic OT
We show in § 3.1 and § 3.2 how to use GENOT to solve OT problems within the same space or across incomparable spaces. On the other hand, numerous real-world problems pose the challenge of the source and target domains being only partially incomparable (Vayer et al., 2018). Therefore, suppose that the source and target space can be decomposed as $\mathcal{X}=\Omega\times\mathcal{\tilde{X}}$ and $\mathcal{Y}=\Omega\times\mathcal{\tilde{Y}}$, respectively. Intuitively, a sample $(\mathbf{u},\mathbf{x})\in\Omega\times\mathcal{\tilde{X}}$ can be interpreted as a structural information $\mathbf{x}$ equipped with a feature $\mathbf{u}$. Afterwards, assume we are given a cost $c:\Omega\times\Omega\rightarrow\mathbb{R}$ to compare features, along with the intra-domain costs $c_{\mathcal{\tilde{X}}},c_{\mathcal{\tilde{Y}}}$. The entropic Fused Gromov-Wasserstein (FGW) problem reads
@@
$\min_{\pi\in\Pi(\mu,\nu)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}Q_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)\,,$ EFGW
@@
where $Q_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\left((\mathbf{u},\mathbf{x}),(\mathbf{v},\mathbf{y}),\mathbf{x}^{\prime},\mathbf{y}^{\prime}\right):=(1-\alpha)\,c(\mathbf{u},\mathbf{v})\,+\,\alpha\,|c_{\mathcal{\tilde{X}}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{\tilde{Y}}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$ and $\alpha\in[0,1]$. This loss combines the pointwise structural distortion and the feature information. The fused problem (EFGW) also admits an unbalanced extension, which can be derived in the same way as (UQEOT) using the quadratic $\mathrm{KL}^{\otimes}$ (Thual et al., 2023).
@@
(U-)GENOT for the Fused Setting. Whether in the balanced or unbalanced setting, we can use our method to learn a specific coupling as soon as it can be estimated from samples. We stress that the discrete solvers we use for (QEOT) and (UQEOT), respectively Peyré et al. (2016)’s and Séjourné et al. (2023b)’s ones, are still applicable in the fused setting. As a result, we can approximate solutions of (EFGW) and its unbalanced counterpart with (U-) GENOT. To illustrate the learning outcome, take a solution $\pi^{\star}_{\alpha}$ of (EFGW). Learning $\pi^{\star}_{\alpha}$ with our method amounts to training vector fields that are conditioned on pairs of modalities from the source domain $v_{t,\theta}(\cdot,|\mathbf{u},\mathbf{x})$, to sample pairs of modalities from the target domain via the flow: $Z\sim\rho$, $\phi_{1}(Z|\mathbf{u},\mathbf{x})=(V,Y)\sim\pi^{\star}_{\alpha}(\cdot|\mathbf{u},\mathbf{x})$. The sampled modalities $(V,Y)$ (i) minimize transport cost quantified by $c$ along the feature (ii) while minimizing the distortion along the structural information quantified by $c_{\tilde{\mathcal{X}}}$ and $c_{\tilde{\mathcal{Y}}}$. In § 5.2, we use fused couplings to enhance the translation between different modalities of cells by adding label information .
Algorithm 1 U-GENOT. Skip teal steps for GENOT.
@@
1:  Require parameters: Batch size $n$; entropic regularization $\varepsilon$; unbalancedness parameter $\tau$; linear, quadratic or fused discrete OT solver $\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}$.
@@
5:      $\mathbf{P}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}\leftarrow\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}\left(\{X_{i}\}_{i=1}^{n},\{Y_{i}\}_{i=1}^{n}\right)\in\mathbb{R}_{+}^{n\times n}$.
@@
10:       ${\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}+\sum_{k}(\eta_{\theta}(X_{k})-n\mathbf{a}_{k})^{2}+(\xi_{\theta}(Y_{k})-n\mathbf{b}_{k})^{2}}$.
@@ 4 Related work and Discussions
Static Neural EOT. While GENOT is the first model to consider the quadratic (and Fused) EOT setting, as well as addressing unbalancedness, various methods have been proposed in the (balanced) linear EOT scenario. The first class of methods solves (LEOT) dual. While some of them (Seguy et al., 2017; Genevay et al., 2019a) do not allow direct sampling from to $\pi^{\star}_{\varepsilon}$aniels et al. (2021) and Mokrov et al. (2023) model the conditional distributions $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. However, these methods might be costly and unstable, as they rely on Langevin sampling during training or inference.
@@
Dynamic Neural EOT. The second class of linear EOT solvers builds upon the link between (LEOT) and the SB problem  (De Bortoli et al., 2021; Chen et al., 2021; Vargas et al., 2021; Gushchin et al., 2022). As simulations are costly in this setting, recent works consider simulation-free training via bridge matching (Shi et al., 2023; Korotin et al., 2023a; Liu et al., 2023; Zhou et al., 2023; Peluchetti, 2023; Tong et al., 2023a, b). While Tong et al. (2023a, b), use mini-batch OT similarly to us, our method is fundamentally different, as we do not build upon the link between EOT and SB. We only use flow matching as a powerful black box to learn a flow from the noise $\rho$ to each conditional $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. Therefore, our approach allows for more flexibility. First (i), the abovementioned methods assume that $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$, since they learn a velocity field (or a drift) by directly bridging $\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathcal{Y}$. Then, they map $\mu$ to $\nu$ with the induced (stochastic) flow. On the other hand, conditionally to each $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, we learn a velocity field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ in the target space, by building paths between noise $\mathbf{z}\in\mathbb{R}^{q}$ and $\mathbf{y}\in\mathcal{Y}\subset\mathbb{R}^{q}$. Then, we map $\rho$ to each $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ with the flows $\phi_{1}(\cdot|\mathbf{x})$, and recover $\nu=\phi_{1}(\cdot|\cdot)\sharp(\rho\otimes\mu)$. Second (ii), since they use discrete couplings to pair the samples $\mathbf{x},\mathbf{y}$ that they bridge, they can only approximate the SB for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. This results from (Bortoli et al., 2023, Prop. 2). In contrast, our method handles any cost, as shown in Fig 2. Third (iii), as they learn a single flow directly transporting $\mu$ to $\nu$, they do not approximate the conditional densities $\pi_{\varepsilon}^{\star}(\mathbf{y}|\mathbf{x})$. Similarly to Tong et al. (2023b), Pooladian et al. (2023) couple samples from $\mu$ and $\nu$, but they only model deterministic maps and assume $\mu=\mathcal{N}(0,I_{d})$. Finally, Bortoli et al. (2023) have recently proposed an augmented bridge matching procedure that preserves couplings. Yet, they still require $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$ and do not approximate conditional densities.
@@
Mini-batches and Biases Quantifying non-asymptotically the bias resulting from minimizing a sample-based GENOT loss, and not its population value, is challenging. The OT curse of dimensionality (Weed & Bach, 2017) has been discussed in generative models relying on mini-batch couplings (Genevay et al., 2019a; Fatras et al., 2021; Tong et al., 2023b, a). Yet, our goal is not to model non-regularized OT, such as a deterministic Monge map, or a Benamou-Brenier vector field. We explicitly target the entropy-regularized OT coupling. Thus, using $\varepsilon\gg 0$ helps to mitigate the curse of dimensionality because of two qualitative factors:
@@
    Statistical. For both linear and quadratic OT, all statistical recovery rates that relate to entropic costs (Genevay et al., 2019b; Zhang et al., 2023) or maps (Pooladian & Niles-Weed, 2021; Rigollet & Stromme, 2022), have a far more favorable regime, with a parametric rate in $\varepsilon>0$ that dodges the curse of dimensionality.
@@
    Computational. While the benefits of employing a large enough $\varepsilon$ in Sinkhorn’s algorithm are widely known, Rioux et al. (2023) have recently shown that as $\varepsilon$ increases, the quadratic OT problem (QEOT) becomes convex, making discrete solvers faster and more reliable.
To demonstrate this aspect, we empirically study the influence of the batch size on GENOT, using data from a recent benchmark (Gushchin et al., 2023), see Fig. LABEL:fig:eot_benchmark_all.
Figure 3: Estimating the  (ULEOT) coupling between Gaussian mixtures with U-GENOT-L ($\varepsilon=0.05$, $\tau_{1}=\tau_{2}=0.98$ and $c=\ell_{2}^{2}$). Left to right: Source (top) $\mu$ and the target $\nu$ (bottom), learned (top) and true (bottom) coupling, learned (top) and true (bottom) re-weighted marginal, and conditional densities. We further showcase the influence of $\tau_{1},\tau_{2}$ in Fig. LABEL:fig:1d_simulated_unbalanced_big.
@@ 5 Experiments
Model rank $\operatorname{B\mathbb{W}_{2}^{2}-UVP}$ rank $\operatorname{cB\mathbb{W}_{2}^{2}-UVP}\downarrow$
LSOT 9.33 9.33
SCONES 7.50 7.67
DiffSB 7.58 7.50
FB-SDE-J 5.12 7.33
FB-SDE-A 6.75 6.75
EgNOT 2.71 5.75
ENOT 6.08 3.00
NOT 4.92 2.92
MLE-SB 3.17 2.67
GENOT-L 1.79 2.08
Table 1: Average ranks of each method across experiments, on estimating the ground-truth EOT coupling $\pi^{\star}_{\varepsilon}$ between each pair of measures $\mu,\nu$, in dimension $d\in\{2,16,64,128\}$ and $\varepsilon\in\{0.1,1,10\}$ of Gushchin et al. (2023)’s benchmark. As them, we measure performances with both $\operatorname{B\mathbb{W}_{2}^{2}-UVP}$ (LABEL:eq:bures_wasserstein) and $\operatorname{cB\mathbb{W}_{2}^{2}-UVP}$ (LABEL:eq:conditional_bures_wasserstein) metrics. See App. LABEL:app:eot_benchmark for details.
@@
Figure 4: Top: Source cell from the early time points (top left) and samples of the conditional distributions of the EOT coupling learned with GENOT for the geodesic cost $d_{\mathcal{M}}$ (middle) and the $\ell_{2}^{2}$ cost (right) projected onto a UMAP (McInnes et al., 2018). Using the geodesic distance as cost function prevents $\pi_{\theta}$ from skipping the intermediate cell population, which is far away from the source cell and terminal cells in $\ell_{2}$ distance. Bottom: UMAP colored according to the uncertainty score $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}^{d_{\mathcal{M}}}(\cdot|X))$ of each source cell $\mathbf{x}$. Target cells are colored in gray.
@@
We show the applicability and versatility of our framework to map within the same space (solving a linear EOT problem) and across incomparable spaces (solving a purely quadratic or fused EOT problem). Metrics and datasets are discussed in App. LABEL:app:metrics and App. LABEL:app:datasets, respectively. Further experimental details or results are reported in App. LABEL:app:experiments. Setups for baselines are listed in App. LABEL:app:competing_methods. Implementation details can be found in App. LABEL:app:implementation. We denote by GENOT-L the GENOT model for solving the linear problem (LEOT), GENOT-Q for the quadratic (QEOT) and GENOT-F for the fused (EFGW) one. When considering the unbalanced counterparts, we add the prefix U. Moreover, when using the conditional mean of a GENOT model, we add the suffix CM. We denote our estimated couplings by $\pi_{\theta}$.
5.1 GENOT for Linear EOT problems

### v4

@@ 3.3 Combining Linear and Quadratic OT
We show in § 3.1 and § 3.2 how to use GENOT to solve OT problems within the same space or across incomparable spaces. On the other hand, numerous real-world problems pose the challenge of the source and target domains being only partially incomparable [88]. Therefore, suppose that the source and target space can be decomposed as $\mathcal{X}=\Omega\times\mathcal{\tilde{X}}$ and $\mathcal{Y}=\Omega\times\mathcal{\tilde{Y}}$, respectively. Intuitively, a sample $(\mathbf{u},\mathbf{x})\in\Omega\times\mathcal{\tilde{X}}$ can be interpreted as a structural information $\mathbf{x}$ equipped with a feature $\mathbf{u}$. Assume we are given a cost $c:\Omega\times\Omega\rightarrow\mathbb{R}$ to compare features, along with the intra-domain costs $c_{\mathcal{\tilde{X}}},c_{\mathcal{\tilde{Y}}}$. The entropic Fused Gromov-Wasserstein (FGW) problem reads
@@
$\min_{\pi\in\Pi(\mu,\nu)}\int_{((\Omega\times\tilde{\mathcal{X}})\times(\Omega\times\tilde{\mathcal{Y}}))^{2}}D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)\,,$ EFGW
@@
where $D_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\left((\mathbf{u},\mathbf{x}),(\mathbf{v},\mathbf{y}),\mathbf{x}^{\prime},\mathbf{y}^{\prime}\right):=(1-\alpha)\,c(\mathbf{u},\mathbf{v})\,+\,\alpha\,|c_{\mathcal{\tilde{X}}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{\tilde{Y}}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$ and $\alpha\in[0,1]$. This loss combines the pointwise structural distortion and the feature information. We refer to the additional cost on the feature information as the fused term. The weighting $\alpha$ allows us to interpolate between purely linear OT on the feature ($\alpha=0$), and purely quadratic OT ($\alpha=1$) on the structural information. Problem (EFGW) also admits an unbalanced extension, derived similarly as (UQEOT) with the quadratic $\mathrm{KL}^{\otimes}$ [82]. An (un)balanced fused EOT coupling always exists (see B.1), it minimizes distortion along the structural information and displacement cost along the features.
@@
(U-)GENOT for the Fused Setting. Whether in the balanced or unbalanced setting, we can use our method to learn a specific coupling as soon as it can be estimated from samples. We stress that the discrete solvers introduced by Peyré et al. [62] and Séjourné et al. [81] we use for (QEOT) and (UQEOT), respectively, are still applicable in the fused setting. As a result, we can approximate solutions of (EFGW) and its unbalanced counterpart with (U-)GENOT. To illustrate the learning outcome, take a solution $\pi^{\star}_{\alpha}$ of (EFGW). Learning $\pi^{\star}_{\alpha}$ with our method amounts to training vector fields that are conditioned on pairs of modalities from the source domain $v_{t,\theta}(\cdot,|\mathbf{u},\mathbf{x})$, to sample pairs of modalities from the target domain via the flow: $Z\sim\rho$, $\phi_{1}(Z|\mathbf{u},\mathbf{x})=(V,Y)\sim\pi^{\star}_{\alpha}(\cdot|\mathbf{u},\mathbf{x})$. The sampled modalities $(V,Y)$ (i) minimize transport cost quantified by $c$ along the feature (ii) while minimizing the distortion along the structural information quantified by $c_{\tilde{\mathcal{X}}}$ and $c_{\tilde{\mathcal{Y}}}$. In § 5.2, we use fused couplings to enhance the translation between different modalities of cells.
Why Does Adding a Fused Term Provide Benefits? Generally, unlike the linear EOT problem (LEOT), the solution $\pi^{\star}_{\varepsilon}$ to the quadratic EOT problem (QEOT) is not unique. This is where adding a fused term helps, as it introduces more constraints. Intuitively, the fused term on additional features $\mathbf{u},\mathbf{v}$ allows us to ’select’ an optimal GW coupling: we trade some GW optimality by choosing a coupling that not only minimizes the distortion of the structural information $\mathbf{x},\mathbf{y}$, but also reduces the displacement cost, quantified by $c$, along the features $\mathbf{u},\mathbf{v}$. Empirically, this significantly mitigates the issue of non-uniqueness in (pure) GW couplings. In § 5.2, we demonstrate that it enhances the stability of our procedure, particularly for single-cell data, as illustrated in Fig. 4 and Fig. LABEL:fig:genot_variance. On a related note, but independently of the addition of the fused term, we investigate in App. LABEL:app:genot_quadratic other strategies to mitigate the non-uniqueness of the GW coupling. Notably, we show that using the geodesic distance on the data manifold as costs, instead of the squared Euclidean distance, helps address this issue on single-cell data. As an alternative, we also introduce an initialization scheme to bias the discrete GW solver to a specific solution based on the neural coupling obtained in the previous iteration (App. LABEL:app:genot_quadratic). Yet, this approach renders GENOT non-simulation free.
3.4 One algorithm to flexibly switch between neural OT problems
@@
1:  Require parameters: Batch size $n$; entropic regularization $\varepsilon$; unbalancedness parameter $\tau$; linear, quadratic or fused discrete OT solver $\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\tau}}$.
@@
5:      $\mathbf{P}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\tau}}\leftarrow\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\tau}}\left(\{X_{i}\}_{i=1}^{n},\{Y_{i}\}_{i=1}^{n}\right)\in\mathbb{R}_{+}^{n\times n}$.
@@
10:       ${\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}+\sum_{k}(\eta_{\theta}(X_{k})-n\mathbf{a}_{k})^{2}+(\xi_{\theta}(Y_{k})-n\mathbf{b}_{k})^{2}}$.
@@
Algorithm 1 U-GENOT. Skip teal steps for GENOT.
Single-cell genomics data is inherently noisy, due to biases incurred by sequencing protocols and the high sparsity of the measurements [34]. Thus, it is indispensable to flexibly switch between configurations of the problem setup. For example, it is often not clear whether the mass conservation constraint should be actually loosened (for example whether to allow for modeling cell death), or whether incorporating prior information for trajectory inference via the quadratic term (e.g. with spatial information [39] or lineage barcoding [42]) is beneficial. Similarly, trying different costs is crucial in single-cell genomics ([17, 35, 39]). The GENOT formulation offers this flexibility by introducing only minor changes into the algorithm. We detail our procedure in Alg. 1, showing the additional steps to switch from GENOT to U-GENOT in teal. Limitations are discussed in § A.
@@ 4 Related work and Discussions
Static Neural EOT. While GENOT is the first model to consider the quadratic (and fused) EOT setting, various methods have been proposed in the linear EOT scenario. The first class of methods solves (LEOT)’s dual. While some of them [73, 27] do not allow direct sampling from to $\pi^{\star}_{\varepsilon}$aniels et al. [15] and Mokrov et al. [55] model $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. However, these methods might be costly and unstable, as they rely on Langevin sampling during training or inference.
@@
Dynamic Neural EOT. The second class of linear EOT solvers builds upon the link between (LEOT) and the SB problem [16, 10, 87, 29]. Although they operate primarily in the balanced setting, Gazdieva et al. [26] recently considered the unbalanced one. As simulations are costly in this setting, recent works consider simulation-free training via bridge matching [76, 41, 46, 96, 61, 84, 85]. While Tong et al. [84, 85] use mini-batch OT, our method is fundamentally different, as we do not build upon the link between EOT and SB. We only use flow matching as a powerful black box to learn a flow from the noise $\rho$ to each conditional $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. Therefore, our approach allows for more flexibility. First (i), the abovementioned methods assume that $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$, since they learn a velocity field (or a drift) by directly bridging $\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathcal{Y}$. Then, they map $\mu$ to $\nu$ with the induced (stochastic) flow. On the other hand, conditionally to each $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, we learn a velocity field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ in the target space, by building paths between noise $\mathbf{z}\in\mathbb{R}^{q}$ and $\mathbf{y}\in\mathcal{Y}\subset\mathbb{R}^{q}$. Then, we map $\rho$ to each $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ with the flows $\phi_{1}(\cdot|\mathbf{x})$, and recover $\nu=\phi_{1}(\cdot|\cdot)\sharp(\rho\otimes\mu)$. Second (ii), they can only approximate the SB for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. This results from [5, Prop. 2]. In contrast, our method handles any cost, as shown in Fig LABEL:fig:coulomb_cost. Third (iii), as they learn a single flow directly transporting $\mu$ to $\nu$, they do not approximate the conditional densities $\pi_{\varepsilon}^{\star}(\mathbf{y}|\mathbf{x})$. Similarly to Tong et al. [85], Pooladian et al. [64] couple samples from $\mu$ and $\nu$, but they only model deterministic maps and assume $\mu=\mathcal{N}(0,I_{d})$. Finally, Bortoli et al. [5] recently proposed an augmented bridge matching procedure that preserves couplings. However, it still requires $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$ and does not approximate conditional densities, limiting its applicability in single-cell genomics tasks.
@@
Mini-batches and Biases Quantifying non-asymptotically the bias resulting from minimizing a sample-based GENOT loss, and not its population value, is challenging. The OT curse of dimensionality [91] has been discussed in generative models relying on mini-batch couplings [27, 22, 85, 84]. Yet, our goal is not to model non-regularized OT, such as a deterministic Monge map, or a Benamou-Brenier vector field. We explicitly target the entropy-regularized OT coupling. Thus, using $\varepsilon\gg 0$ helps to mitigate the curse of dimensionality because of two qualitative factors:
@@
    Statistical. For both linear and quadratic OT, all statistical recovery rates that relate to entropic costs [28, 95], maps [63, 66], or couplings [90], have a far more favorable regime, with a parametric rate in $\varepsilon>0$ that dodges the curse of dimensionality.
@@
    Computational. While the benefits of employing a large enough $\varepsilon$ in Sinkhorn’s algorithm are widely known, Rioux et al. [67] have recently shown that as $\varepsilon$ increases, the quadratic OT problem (QEOT) becomes convex, making discrete solvers faster and more reliable.
To demonstrate this aspect, we empirically study the influence of the batch size on GENOT, using a recent benchmark [30], with known true linear EOT coupling $\pi_{\varepsilon}^{\star}$, see Fig. LABEL:fig:eot_benchmark_all.
@@ 5 Experiments
Figure 2: Left: Source cell from the early time points (top left) and samples of the conditional distributions of the EOT coupling learned with GENOT for the geodesic cost $d_{\mathcal{M}}$ (middle) and the $\ell_{2}^{2}$ cost (right) projected onto a UMAP [53], along with biological assessment of the learnt dynamics (TSI score, CT error § LABEL:app:metrics_single_cell, Fig. LABEL:fig:pancreas_transitions). Right: UMAP colored according to the uncertainty score $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}^{d_{\mathcal{M}}}(\cdot|X))$ of each source cell $\mathbf{x}$. Target cells are colored in gray.
@@
While there is no evidence that cells exactly evolve according to an entropic OT plan, leveraging OT is an established way to realign cells. We also evaluate the performance on simulated data in settings which closed-form solutions of the EOT coupling are available for. Metrics and datasets are discussed in App. LABEL:app:metrics and App. LABEL:app:datasets, respectively. Further experimental details or results are reported in App. LABEL:app:experiments. Setups for baselines are listed in App. LABEL:app:competing_methods. Implementation details can be found in App. LABEL:app:implementation. We denote by GENOT-L the GENOT model for solving the linear problem (LEOT), GENOT-Q for the quadratic (QEOT) and GENOT-F for the fused (EFGW) one. When considering the unbalanced counterparts, we add the prefix U. Moreover, when using the conditional mean of a GENOT model, we add the suffix CM. We denote our learned neural couplings by $\pi_{\theta}$.
5.1 GENOT-L for modeling cell trajectories